OpenStudy (anonymous):

The revenue function for a production by a theatre group is R(t) = -50t^2 + 300t where t is the ticket price in dollars. The cost function for the production is C(t) = 600-50t. Determine the ticket price that will allow the production to break even.

OpenStudy (anonymous):

The answer is $3.00 and I got after equating the two equations and using quadratic formula to find x. x = 4, x = 3. I'm confused!

OpenStudy (anonymous):

Just set R(t) = C(t) and solve for t -50t^2+300t = 600-50t -50t^2+350t -600 = 0 Now solve for T using the quadratic formula.

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

yeah i get two solutions as well maybe they were not paying attention when they made up the problem

OpenStudy (anonymous):

Looks like you can have two different answers. Here's a visual if anyone's interested.

OpenStudy (anonymous):

Paste this into google if you can't view the attachment: -50t^2 + 300t and 600-50t